OpenStudy (anonymous):

how do you solve -3u=u^2

geerky42 (geerky42):

Try and divide both side by u.

OpenStudy (anonymous):

So then u= -3?

geerky42 (geerky42):

Yeah, but you are missing something, sorry, I jumped to conclusion. Do it a different way. add both sides by 3u and we have \(0 = u^2+3u\) Factor u out: \(0 = u(u+3)\) So \(\boxed{u=0,-3}\)

geerky42 (geerky42):

does that make sense?

OpenStudy (anonymous):

oh that makes more sense thank you ^_^